OpenStudy (anonymous):

3 ____ 7 ____2 ____4 = 13 anyone know what this is?

Directrix (directrix):

@LeonSquishy What are the instructions for the problem?

Directrix (directrix):

Maybe, fill in the operation(s) that make the number sentence true?

OpenStudy (anonymous):

Use the operations symbols +, - and x to make the following equation true.

OpenStudy (anonymous):

+, x, - remember, you have to use the order of operations so mult & div would come first, in order from left to right followed by + or - in order from left to right.

OpenStudy (anonymous):

3+7 x 2 - 4 = 3 + 14 - 4 = 17 - 4 = 13

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

you're welcome!